# Semiclassical descriptions of dissipative dynamics of strongly interacting Bose gases in optical lattices

### v2

@@ Abstract
We develop semiclassical methods for describing real-time dynamics of dissipative Bose-Hubbard systems in a strongly interacting regime, which can be realized in experiments with ultracold gases. Specifically, we present two kinds of SU(3) truncated Wigner approximation (TWA) for solving an effective Lindblad master equation of the systems. The first one employs the discrete TWA for finite levels or spin systems and is based on a classical equation of motion in which the onsite dissipation term, as well as the onsite interaction term, is linearized in the phase space variables. The other approach utilizes a stochastic Langevin equation, including decoherence effects in terms of nonlinear drift force and stochastic force terms, in which the initial conditions of trajectories are weighted with a quasiprobability distribution for a typical initial quantum state. We apply these methods to the systems with two-body losses and compare their results with the exact numerical solutions for a small system. We show that the former approach can simulate correctly longer-time dynamics than the latter one. We also calculate the time evolution for a large size setup that is comparable to experiments. We numerically demonstrate that the discrete TWA approach is able to qualitatively capture the continuous quantum Zeno effect on dynamics subjected to a gradual change of the ratio between the hopping amplitude and the onsite interaction across the superfluid-Mott insulator crossover, which has been observed experimentally.
@@ I Introduction
For higher dimensional open systems, there are various time-dependent mean-field approaches that can approximately evaluate the time dependence of local observables governed by microscopic master equations for mixed states. For instance, in Ref. Tomita et al. (2017), the time-dependent Gutzwiller approximation method Diehl et al. (2010); Tomadin et al. (2011) is employed to analyze the time evolution of a dissipative Bose-Hubbard system with two-body losses. It has been numerically shown that a site decoupling ansatz for many-body mixed states can nicely recover some qualitative features of the Zeno effect that can be observed in the local atomic density Tomita et al. (2017). However, in the Gutzwiller approach, for inducing crossover from a Mott insulator phase to a superfluid phase, one needs to introduce artificial noises as small perturbations. Such noises inevitably cause uncontrolled errors, leading to quantitative inaccuracy in the results. Moreover, due to its site decoupling nature, the Gutzwiller approach cannot properly describe offsite correlation functions. This drawback significantly limits the application of the Gutzwiller approach to ultracold-gas systems, e.g., not being able to evaluate the momentum distribution of atomic gases, an indispensable observable measured via time-of-flight images, because it is usually computed as a sum of offsite single-particle correlation functions in real space.
@@
In a typical way of semiclassically expressing the dynamics of the isolated Bose-Hubbard model in its weakly interacting regime, trajectories are generated by the Hamilton equation for complex-valued field amplitudes, i.e., the discrete Gross-Pitaevskii (GP) equation Blakie et al. (2008); Polkovnikov (2010); Nagao et al. (2019); Ozaki et al. (2020). Although the TWA based on the GP equation no longer provides an appropriate description in the strongly interacting regime, there is an alternative representation, namely, the SU(3) TWA, which exactly describes onsite particle and hole fluctuations in strongly interacting states Davidson and Polkovnikov (2015); Fujimoto et al. (2020); Nagao et al. (2021). In this method, the equation of motion in the classical limit is given as the Hamilton equation for SU(3) pseudospin variables, and it is formally equivalent to the time dependent variational equation in the single-site Gutzwiller mean-field theory Nagao et al. (2021). However, in the case of a driven dissipative system, which is not governed by the von-Neumann equation for isolated systems, the classical trajectories cannot be accurately reproduced via the Hamilton equations alone. This is because dissipative forces emerge in addition to the Hamiltonian dynamics, representing the effects of decoherence that influence the coherent time evolution.
@@
We demonstrate that these two different TWA approaches can qualitatively describe the continuous quantum Zeno effect in the non-unitary dynamics with strong dissipation. However, the resulting threshold of the dissipation strength for exhibiting the Zeno effect is different, depending on which approach is employed for the simulation. In particular, we show that, because of the linearity of the non-unitary contributions in the classical limit, the dTWA representation can produce reasonable threshold values of the dissipation strength separating Zeno and non-Zeno regimes in comparison with the experimental results on the time evolution of the atomic density as well as the momentum distribution.
@@ II Model
Figure 1: Schematic illustration of the analog quantum simulation experiment for a driven dissipative Bose-Hubbard system in Ref. Tomita et al. (2017). In order to effectively induce two-body loss dissipation, a PA laser is shined to an atomic system comprised of ${}^{174}{\rm Yb}$ atoms in a three-dimensional optical lattice.
@@
In Eq. (1), the dissipative contribution in the presence of the PA laser is described by the non-unitary term $L_{2}[{\hat{\rho}}_{\rm eff}]$. For the setup that we consider here, the onsite losses of atoms are represented by
@@ III.1 General remarks
Before proceeding, we briefly make key remarks on the SU(${\cal N}$) TWA method that we employ. This method is a phase-space approach for numerically simulating real-time dynamics of interacting spin systems for spin $S\geq 1/2$ Wurtz et al. (2018). Within the approximation, the time evolution of a quantum spin model reduces to an ensemble of fluctuating trajectories on the phase space of classical SU(${\cal N}$) spins. The stochastic nature arises form the presence of quantum fluctuations in the initial state. The idea of the SU(${\cal N}$) TWA relies on the fact that any local non-linear operator in spin-$S$ systems can be mapped to a linear combination of generators of the SU(${\cal N}$) group, where ${\cal N}=2S+1$. For instance, for an $S=1$ magnet with single-ion uniaxial anisotropy, its quadratic interaction term, i.e., $\sum_{i}({\hat{S}}^{z}_{i})^{2}$, is transformed into a linear term of generators of the SU(3) group. The TWA applied to such a linearized version of the model significantly improves beyond the capability of the naive approximation for the original model written in terms of the SU(2) spin operators, i.e., ${\hat{S}}^{x}_{i}$, ${\hat{S}}^{y}_{i}$, and ${\hat{S}}^{z}_{i}$ Davidson and Polkovnikov (2015). In previous studies, the SU(${\cal N}$) TWA has been applied to experiments for magnetic atoms to realize a large-$S$ spin system Lepoutre et al. (2019), spin-1 systems with single-ion uniaxial anisotropy and strongly interacting Bose-Hubbard systems Davidson and Polkovnikov (2015); Fujimoto et al. (2020); Nagao et al. (2021), and cluster representations of spin-$1/2$ systems in one dimension Wurtz et al. (2018).
@@ III.2 SU(3) matrix representation of the dissipative Bose-Hubbard system
To apply the SU(${\cal N}$) TWA to the effective Lindblad master equation in Eq. (1), we derive a reduced model in a projected Hilbert space spanned by relevant Fock states. In the experiment, the onsite interaction $U$ is typically tuned to be in sufficiently large compared to the hopping amplitude $J$, i.e., $U\gg J$. This condition implies that the relevant single-particle states of onsite occupation are only a few states around a mean filling state Altman and Auerbach (2002); Nagao et al. (2018, 2021). In this paper, we assume that the mean filling state at each site is the unit filling state $|1\rangle$, and the relevant onsite excitations around it are the vacuum state $|0\rangle$ and the doubly occupied state $|2\rangle$, therefore with ${\cal N}=3$. In such a reduced space, written as $\bigotimes_{j}\{|0\rangle,|1\rangle,|2\rangle\}_{j}$, any physical operator, including the Hamiltonian operator and the effective density matrix, can be represented as an expansion in tensor products of three-dimensional SU(3) base matrices such as the Gell-Mann matrices Georgi (2000).
@@ III.3 Method I: dTWA
Let us now explain the dTWA Schachenmayer et al. (2015); Piñeiro Orioli et al. (2017); Kunimi et al. (2021); Singh and Weimer (2022) to approximately calculate the time evolution of the dissipative interacting SU(3) spin system. The dTWA method for finite levels or spin systems formally uses a phase point operator to expand the density matrix operator. For the case of SU(3) spin systems with $N$ sites, an appropriate definition of phase point operator is given via its matrix representation, i.e., ${\mathscr{A}}^{\otimes N}=\bigotimes_{j=1}^{N}{\mathscr{A}}^{[j]}$ with ${\mathscr{A}}^{[j]}=\frac{1}{3}I+\frac{1}{2}{\bm{r}}^{(j)}\cdot{\bm{T}}$ Nagao et al. (2021) and the normalization conditions of $T_{\alpha}$ being ${\rm Tr}(T_{\alpha}T_{\beta})=2\delta_{\alpha\beta}$. Moreover, ${\bm{r}}^{(j)}=(r^{(j)}_{1},\cdots,r^{(j)}_{8})$ represents a spinor vector of eight real-valued discrete parameters. Using the phase point operator, one can expand the initial density matrix at $t=0$ such that $\rho(t=0)=\prod_{j}\sum_{{\bm{r}}^{(j)}}{\mathscr{W}}({\bm{r}}^{(1)},\cdots,{\bm{r}}^{(N)}){\mathscr{A}}^{\otimes N}$. The expansion coefficient ${\mathscr{W}}({\bm{r}}^{(1)},\cdots,{\bm{r}}^{(N)})$ is interpreted as a discrete Wigner function associated with the phase point operator. For a direct product state over real space, the discrete Wigner function is given as a spatially factorized distribution function, i.e., ${\mathscr{W}}({\bm{r}}^{(1)},\cdots,{\bm{r}}^{(N)})=\prod_{j}{\mathscr{W}}_{1}^{[j]}({\bm{r}}^{(j)})$. The local distribution ${\mathscr{W}}_{1}^{[j]}({\bm{r}}^{(j)})$ depends only on ${\bm{r}}^{(j)}$. The explicit form of ${\mathscr{W}}_{1}^{[j]}({\bm{r}}^{(j)})$ used in the main calculations is provided in Appendix D.
With the initial density matrix expanded with the phase point operator described above, we now consider the non-unitary time evolution of the phase point operator, which is also generated by the effective Lindblad master equation in Eq. (1). Let us introduce a time-dependent phase point operator at time $t$ as ${\hat{\mathscr{A}}}^{\otimes N}_{t}[\{{\bm{r}}^{(k)}\}_{k=1}^{N}]$ with the initial condition ${\hat{\mathscr{A}}}^{\otimes N}_{t=0}={\hat{\mathscr{A}}}^{\otimes N}$. This operator has to evolve according to
@@ III.4 Method II: Fokker-Planck TWA
where $\alpha=1,2,\cdots,8$ and $q=1,2$. The drift term is translated into the deterministic forces $f_{\alpha}^{(j)}dt=(f_{{\rm unitary},\alpha}^{(j)}+f_{{\rm PA},\alpha}^{(j)})dt$, including non-unitary contributions $f_{{\rm PA},\alpha}^{(j)}$ from the PA laser, while the diffusion term gives rise to the stochastic forces $\sum_{q=1}^{2}g_{\alpha,q}^{(j)}\cdot dw_{q}^{(j)}$. In the limit of $\Gamma_{\rm PA}=0$, Eq. (14) reduces to the Hamilton equation for SU(3) variables Nagao et al. (2021), i.e., $dx_{\alpha}^{(j)}=-(i/\hbar)\{x_{\alpha}^{(j)},H_{\rm BH}^{W}\}_{\rm P.B.}dt=f_{{\rm unitary},\alpha}^{(j)}dt$. The local noises $dw_{q}^{(j)}={\cal O}(dt^{1/2})$ must obey a Gaussian probability distribution around the zero mean value, whose variance is specified by the normalization conditions $dw_{q}^{(j)}dw_{q^{\prime}}^{(j^{\prime})}=dt\delta_{q,q^{\prime}}\delta_{j,j^{\prime}}$. In addition, the dot product between $g_{\alpha,q}^{(j)}$ and $dw_{q}^{(j)}$ has to be interpreted as the Itô product for random variables Gardiner (2009). The analytical expressions of $f_{{\rm PA},\alpha}^{(j)}$ and $g_{\alpha,q}^{(j)}$ are provided in Eqs. (E) and (30), respectively, in Appendix E, which can recover the original FP equation. In particular, the results of $g_{\alpha,q}^{(j)}$ are consistent with the factorization property of the diffusion matrix, i.e., $D^{(j)}=B_{j}B^{\rm T}_{j}$, where $B_{j}=[{\bm{g}}_{q=1}^{(j)},{\bm{g}}_{q=2}^{(j)},{\bm{0}},\cdots,{\bm{0}}]$ is an $8\times 8$ matrix.
The force vectors arising in the presence of the PA laser are given as nonlinear functions of $x_{\alpha}^{(j)}$, and the nonlinearity is measured by $\Gamma_{\rm PA}$. The FP-TWA for the Lindblad master equation is therefore not exact even in the zero hopping limit $J=0$, implying that the valid timescale of the approximation becomes shorter as $\Gamma_{\rm PA}$ increases. We will numerically compare the results obtained by the FP-TWA with those obtained by the dTWA to examine the limitations of the FP-TWA in detail in Sec. IV.4. When $\Gamma_{\rm PA}=0$ and $J=0$, the Langevin equation reduces to a linear equation, and the FP-TWA becomes exact, as expected. Therefore, for dissipative systems, these two approaches, i.e., the Bopp operator method and the phase point operator method, lead to noticeable differences in the presence of non-unitary contributions.

### v3

# Discrete-phase-space method for driven-dissipative dynamics of strongly interacting bosons in optical lattices

@@ Abstract
We develop a discrete truncated Wigner method to analyze the real-time evolution of dissipative SU(${\cal N}$) spin systems coupled with a Markovian environment. This semiclassical approach is not only numerically efficient but also particularly capable of accurately capturing local loss processes due to its local linearity in the dynamical equations. We apply the method to a state-of-the-art experiment involving an analog quantum simulator of a three-dimensional dissipative Bose-Hubbard model in a strongly interacting regime. Our numerical results show good agreement with experimental data, specifically capturing the continuous quantum Zeno effect in the dynamics subjected to a gradual change of the ratio between the hopping amplitude and the onsite interaction across the superfluid-Mott insulator crossover. Furthermore, we present comparative analyses with the continuous truncated Wigner method, derived as an effective Fokker-Planck equation for SU(${\cal N}$) classical spin variables, showing that the discrete method outperforms the continuous one in simulating the long-time dynamics of SU(2) and SU(3) spin models. The discrete phase space framework offers a versatile and powerful tool for exploring a wide range of open quantum many-body systems in dimensions higher than one dimension, where numerically exact methods are impractical due to the exponential growth of the Hilbert space dimension.
@@ I Introduction
For higher-dimensional open systems, there are various time-dependent mean-field approaches that can approximately evaluate the time dependence of local observables governed by microscopic master equations for mixed states. For instance, in Ref. Tomita et al. (2017), the time-dependent Gutzwiller approximation method Diehl et al. (2010); Tomadin et al. (2011) is employed to analyze the time evolution of a dissipative Bose-Hubbard system with two-body losses. It has been numerically shown that a site decoupling ansatz for many-body mixed states can nicely recover some qualitative features of the Zeno effect that can be observed in the local atomic density Tomita et al. (2017). However, in the Gutzwiller approach, for inducing crossover from a Mott insulator phase to a superfluid phase, one needs to introduce artificial noises as small perturbations. Such noises inevitably cause uncontrolled errors, leading to quantitative inaccuracy in the results. Moreover, due to its site decoupling nature, the Gutzwiller approach cannot properly describe offsite correlation functions. This drawback significantly limits the application of the Gutzwiller approach to ultracold-gas systems, e.g., not being able to evaluate the momentum distribution of atomic gases, an indispensable observable measured via time-of-flight images, because it is usually computed as a sum of offsite single-particle correlation functions in real space.
@@
In a typical way of semiclassically expressing the dynamics of the isolated Bose-Hubbard model in its weakly interacting regime, trajectories are generated by the Hamilton equation for complex-valued field amplitudes, i.e., the discrete Gross-Pitaevskii (GP) equation Blakie et al. (2008); Polkovnikov (2010); Nagao et al. (2019); Ozaki et al. (2020). Although the TWA based on the GP equation no longer provides an appropriate description in the strongly interacting regime, there is an alternative representation, namely, the SU(${\cal N}$) TWA, which exactly describes onsite particle and hole fluctuations in strongly interacting states Davidson and Polkovnikov (2015); Fujimoto et al. (2020); Nagao et al. (2021). In this method, the equation of motion in the classical limit is given as the Hamilton equation for SU(${\cal N}$) pseudospin variables, and it is formally equivalent to the time dependent variational equation in the single-site Gutzwiller mean-field theory Nagao et al. (2021). However, in the case of a driven dissipative system, which is not governed by the von-Neumann equation for isolated systems, the classical trajectories cannot be accurately reproduced via the Hamilton equations alone. This is because dissipative forces emerge in addition to the Hamiltonian dynamics, representing the effects of decoherence that influence the coherent time evolution.
@@
We demonstrate that these two different TWA approaches can qualitatively describe the continuous quantum Zeno effect in the nonunitary dynamics with strong dissipation. However, the resulting threshold of the dissipation strength for exhibiting the Zeno effect is different, depending on which approach is employed for the simulation. In particular, we show that, because of the linearity of the nonunitary contributions in the classical limit, the dTWA representation can produce reasonable threshold values of the dissipation strength separating Zeno and non-Zeno regimes in comparison with the experimental results on the time evolution of the atomic density as well as the momentum distribution. The local linearity of the dTWA allows it not only to predict reasonable values of the Zeno threshold but also to provide valuable insights into the qualitative behavior of long-time loss dynamics. Indeed, the dTWA approach can capture second-order perturbation processes of hopping in the original Bose-Hubbard representation, as manifested by an emergent scaling law for atomic density decay after rescaling with an appropriate time unit. Additionally, we show that the dTWA approach can qualitatively reproduce the power-law decay in the long-time evolution of atomic density, with the extracted exponents exhibiting a dependence on the dissipation strength similar to the exact results. These features cannot be described by the FP-TWA approach.
@@ II Model
Figure 1: Schematic illustration of the analog quantum simulation experiment for a driven-dissipative Bose-Hubbard system in Ref. Tomita et al. (2017). In order to effectively induce two-body loss dissipation, a PA laser is shined to an atomic system comprised of ${}^{174}{\rm Yb}$ atoms in a three-dimensional optical lattice.
@@
In Eq. (1), the dissipative contribution in the presence of the PA laser is described by the nonunitary term $L_{2}[{\hat{\rho}}_{\rm eff}]$. For the setup that we consider here, the onsite losses of atoms are represented by
@@ III.1 General remarks
Before proceeding, we briefly make key remarks on the SU(${\cal N}$) TWA method that we employ. This method is a phase-space approach for numerically simulating real-time dynamics of interacting spin systems for spin $S\geq 1/2$ Wurtz et al. (2018). Within the approximation, the time evolution of a quantum spin model reduces to an ensemble of fluctuating trajectories on the phase space of classical SU(${\cal N}$) spins. The stochastic nature arises from the presence of quantum fluctuations in the initial state. The idea of the SU(${\cal N}$) TWA relies on the fact that any local non-linear operator in spin-$S$ systems can be mapped to a linear combination of generators of the SU(${\cal N}$) group, where ${\cal N}=2S+1$. For instance, for an $S=1$ magnet with single-ion uniaxial anisotropy, its quadratic interaction term, i.e., $\sum_{i}({\hat{S}}^{z}_{i})^{2}$, is transformed into a linear term of generators of the SU(3) group. The TWA applied to such a linearized version of the model significantly improves beyond the capability of the naive approximation for the original model written in terms of the SU(2) spin operators, i.e., ${\hat{S}}^{x}_{i}$, ${\hat{S}}^{y}_{i}$, and ${\hat{S}}^{z}_{i}$ Davidson and Polkovnikov (2015). In previous studies, the SU(${\cal N}$) TWA has been applied to experiments for magnetic atoms to realize a large-$S$ spin system Lepoutre et al. (2019), spin-1 systems with single-ion uniaxial anisotropy and strongly interacting Bose-Hubbard systems Davidson and Polkovnikov (2015); Fujimoto et al. (2020); Nagao et al. (2021), and cluster representations of spin-$1/2$ systems in one dimension Wurtz et al. (2018).
@@ III.2 SU(3) matrix representation of the dissipative Bose-Hubbard system
To apply the SU(${\cal N}$) TWA to the effective Lindblad master equation in Eq. (1), we derive a reduced model in a projected Hilbert space spanned by relevant Fock states. In the experiment, the onsite interaction $U$ is typically tuned to be sufficiently large compared to the hopping amplitude $J$, i.e., $U\gg J$. This condition implies that the relevant single-particle states of onsite occupation are only a few states around a mean filling state Altman and Auerbach (2002); Nagao et al. (2018, 2021). In this paper, we assume that the mean filling state at each site is the unit filling state $|1\rangle$, and the relevant onsite excitations around it are the vacuum state $|0\rangle$ and the doubly occupied state $|2\rangle$, therefore with ${\cal N}=3$. In such a reduced space, written as $\bigotimes_{j}\{|0\rangle,|1\rangle,|2\rangle\}_{j}$, any physical operator, including the Hamiltonian operator and the effective density matrix, can be represented as an expansion in tensor products of three-dimensional SU(3) base matrices such as the Gell-Mann matrices Georgi (2000).
@@ III.3 Method I: dTWA
Let us now explain the dTWA Schachenmayer et al. (2015); Piñeiro Orioli et al. (2017); Kunimi et al. (2021); Singh and Weimer (2022) to approximately calculate the time evolution of the dissipative interacting SU(3) spin system. The dTWA method for finite levels or spin systems formally uses a phase point operator to expand the density matrix operator. For the case of SU(3) spin systems with $N$ sites, an appropriate definition of phase point operator is given via its matrix representation, i.e., ${\mathscr{A}}^{\otimes N}=\bigotimes_{j=1}^{N}{\mathscr{A}}^{[j]}$ with ${\mathscr{A}}^{[j]}=\frac{1}{3}I+\frac{1}{2}{\bm{r}}^{(j)}\cdot{\bm{T}}$ Nagao et al. (2021) and the normalization conditions of $T_{\alpha}$ being ${\rm Tr}(T_{\alpha}T_{\beta})=2\delta_{\alpha\beta}$. Moreover, ${\bm{r}}^{(j)}=(r^{(j)}_{1},\cdots,r^{(j)}_{8})$ represents a spinor vector of eight real-valued discrete parameters. Using the phase point operator, one can expand the initial density matrix at $t=0$ such that $\rho(t=0)=\prod_{j}\sum_{{\bm{r}}^{(j)}}{\mathscr{W}}({\bm{r}}^{(1)},\cdots,{\bm{r}}^{(N)}){\mathscr{A}}^{\otimes N}[{\bm{r}}^{(1)},\cdots,{\bm{r}}^{(N)}]$. The expansion coefficient ${\mathscr{W}}({\bm{r}}^{(1)},\cdots,{\bm{r}}^{(N)})$ is interpreted as a discrete Wigner function associated with the phase point operator. For a direct product state over real space, the discrete Wigner function is given as a spatially factorized distribution function, i.e., ${\mathscr{W}}({\bm{r}}^{(1)},\cdots,{\bm{r}}^{(N)})=\prod_{j}{\mathscr{W}}_{1}^{[j]}({\bm{r}}^{(j)})$. The local distribution ${\mathscr{W}}_{1}^{[j]}({\bm{r}}^{(j)})$ depends only on ${\bm{r}}^{(j)}$. The explicit form of ${\mathscr{W}}_{1}^{[j]}({\bm{r}}^{(j)})$ used in the main calculations is provided in Appendix D.
With the initial density matrix expanded with the phase point operator described above, we now consider the nonunitary time evolution of the phase point operator, which is also generated by the effective Lindblad master equation in Eq. (1). Let us introduce a time-dependent phase point operator at time $t$ as ${\hat{\mathscr{A}}}^{\otimes N}_{t}[\{{\bm{r}}^{(k)}\}_{k=1}^{N}]$ with the initial condition ${\hat{\mathscr{A}}}^{\otimes N}_{t=0}={\hat{\mathscr{A}}}^{\otimes N}$. This operator has to evolve according to
@@ III.4 Method II: Fokker-Planck TWA
where $\alpha=1,2,\cdots,8$ and $q=1,2$. The drift term is translated into the deterministic forces $f_{\alpha}^{(j)}dt=(f_{{\rm unitary},\alpha}^{(j)}+f_{{\rm PA},\alpha}^{(j)})dt$, including nonunitary contributions $f_{{\rm PA},\alpha}^{(j)}$ from the PA laser, while the diffusion term gives rise to the stochastic forces $\sum_{q=1}^{2}g_{\alpha,q}^{(j)}\cdot dw_{q}^{(j)}$. In the limit of $\Gamma_{\rm PA}=0$, Eq. (14) reduces to the Hamilton equation for SU(3) variables Nagao et al. (2021), i.e., $dx_{\alpha}^{(j)}=-(i/\hbar)\{x_{\alpha}^{(j)},H_{\rm BH}^{W}\}_{\rm P.B.}dt=f_{{\rm unitary},\alpha}^{(j)}dt$. The local noises $dw_{q}^{(j)}={\cal O}(dt^{1/2})$ must obey a Gaussian probability distribution around the zero mean value, whose variance is specified by the normalization conditions $dw_{q}^{(j)}dw_{q^{\prime}}^{(j^{\prime})}=dt\delta_{q,q^{\prime}}\delta_{j,j^{\prime}}$. In addition, the dot product between $g_{\alpha,q}^{(j)}$ and $dw_{q}^{(j)}$ has to be interpreted as the Itô product for random variables Gardiner (2009). The analytical expressions of $f_{{\rm PA},\alpha}^{(j)}$ and $g_{\alpha,q}^{(j)}$ are provided in Eqs. (E) and (30), respectively, in Appendix E, which can recover the original FP equation. In particular, the results of $g_{\alpha,q}^{(j)}$ are consistent with the factorization property of the diffusion matrix, i.e., $D^{(j)}=B_{j}B^{\rm T}_{j}$, where $B_{j}=[{\bm{g}}_{q=1}^{(j)},{\bm{g}}_{q=2}^{(j)},{\bm{0}},\cdots,{\bm{0}}]$ is an $8\times 8$ matrix.
The force vectors arising in the presence of the PA laser are given as nonlinear functions of $x_{\alpha}^{(j)}$, and the nonlinearity is measured by $\Gamma_{\rm PA}$. The FP-TWA for the Lindblad master equation is therefore not exact even in the zero hopping limit $J=0$, implying that the valid timescale of the approximation becomes shorter as $\Gamma_{\rm PA}$ increases. We will numerically compare the results obtained by the FP-TWA with those obtained by the dTWA to examine the limitations of the FP-TWA in detail in Sec. IV.4. When $\Gamma_{\rm PA}=0$ and $J=0$, the Langevin equation reduces to a linear equation, and the FP-TWA becomes exact, as expected. Therefore, for dissipative systems, these two approaches, i.e., the Bopp operator method and the phase point operator method, lead to noticeable differences in the presence of nonunitary contributions.